OpenStudy (anonymous):

On the same graph make accurate sketches of the functions y= 2x^2 -2 and y= 3x+12 . Clearly indicate all intercepts with axes or functions. How to solve this? thank you!!

OpenStudy (anonymous):

For y=2x^2 -2 : Start with x=0, then y= -2. To find the x-intercepts, let y=0. Then solve for x. You'll get x=1, -1. Try putting in other random, but small points on the graph. For y=3x+12. Start with x= 0, then you get y= 12, which is the y-intercept. To get the x intercept, let y = 0, you get x=-4 for the intercept. Connect those points together and you've got your graph!

OpenStudy (anonymous):

y-12 over 3=x plug that into y=2x^2-2 and work it out i hope thats how u do it if i recall it is that correct meganch iu

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

ok.. I'm doing the questions right now..

OpenStudy (shadowfiend):

Don't forget to give a medal if someone was helpful!